I need help with this problem
applications of quadratic equation
find two numbers whose sum is 7 and whose product is 10.

Answer by checkley71(8403)

X+Y=7
X*Y=10 THEN
X=7-Y AND SUBSTITUTING (7-Y)FOR X IN THE EQUATION 2
WE GET
(Y-7)Y=10
Y^2-7Y=10
Y^2-7Y-10=0
(Y-5)(Y-2)=0
Y-5=0
Y=5 SOLUTION
Y-2=0
Y=2 SOLUTION
PROOF
5+2=7
7=7
2*5=10
10=10